OpenStudy (anonymous):

find the equation of circle passing through the points (1,5)(5,3)&(3,-1).

OpenStudy (amistre64):

those form a triangle that i believe you would need to find the medians of, and the center of the circle is where they intersect

OpenStudy (amistre64):

or you could try something that might be a little tricker and find a solution to equating all these to the same distance from a point

OpenStudy (amistre64):

sqrt((x-1)^2+(y-5)^2) sqrt((x-5)^2+(y-3)^2) sqrt((x-3)^2+(y+1)^2) and see if you can equate them :)

OpenStudy (amistre64):

sqrt((x-1)^2+(y-5)^2) = sqrt((x-5)^2+(y-3)^2) sqrt((x-5)^2+(y-3)^2) = sqrt((x-3)^2+(y+1)^2) (x-1)^2+(y-5)^2 = (x-5)^2+(y-3)^2 (x-5)^2+(y-3)^2 = (x-3)^2+(y+1)^2 x^2 -2x + 1+ y^2 -10y + 25 = x^2 -10x + 25 + y^2 -6y + 9 x^2 -10x + 25 + y^2 -6y + 9 = x^2 -6x + 9 + y^2 +2y +1 -2x -10y + 26 = -10x -6y + 34 -10x -6y + 34 = -6x +2y +10

OpenStudy (amistre64):

8x -4y = 8 -4x -8y = -24

OpenStudy (amistre64):

now that should be easier for you to determine x and y with

OpenStudy (amistre64):

and with any luck, i didnt mess it up any lol

OpenStudy (anonymous):

amistre64 how about i apply \[x^2+y^2+2gx+2fy+c=0\] on each of them

OpenStudy (amistre64):

give it a shot, it might work out

OpenStudy (anonymous):

k i did and got three equations, what do i do next

OpenStudy (amistre64):

how many unknowns? or rather, what equations did you come up with?

OpenStudy (anonymous):

2g+10f+c=-26, 10g+6f+c=-34 and 6g-2f+c=-10

OpenStudy (anonymous):

i guess i'll have to find g,f and c.

OpenStudy (amistre64):

3 eqs in 3 unknowns is doable ...

OpenStudy (amistre64):

my way got me (2,2) as a center, and if i plug in x and y into the distances i get: (1)^2+(-3)^2 = (-3)^2+(1)^2 (-3)^2+(-1)^2 = (-1)^2+(3)^2 d^2 = 10 (x-2)^2 + (y-2)^2 = 10 might be it

OpenStudy (amistre64):

if the given points fit as solutions; its good :)

OpenStudy (anonymous):

dont u think i solve them simultaneously?